OpenStudy (anonymous):

How do I check if this solution is extraneous? Please help. sqrt{2x+6}-sqrt{x-1}=2

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x-1}=2\]

OpenStudy (kc_kennylau):

Since it's the second time you ask this kind of questions, I'm gonna send you a video to help you: http://www.youtube.com/watch?v=711pdW8TbbY

OpenStudy (kc_kennylau):

Have you tried solving it? :)

OpenStudy (kc_kennylau):

OS values the learning process rather than the answer. Please work out the solution by yourself and we will guide you.

OpenStudy (tkhunny):

Is \(2x+6 \ge 0\)? Is \(x-1 \ge 0\)?